(x/x^2-4)-1(x+2)=2
Answer by edjones(8007)

(x/(x^2-4))-1/(x+2)=2 Assuming that this is the problem.
x/((x+2)(x-2)) - 1/(x+2)=2
x-(x-2)=2(x^2-4) multiply each side by (x-2)(x+2)
x-x+2=2x^2-8
2x^2-10=0
2(x^2-5)=0
x^2=5
x=sqrt(5), x=-sqrt(5)
.
Ed